There are 1000 students in the senior class at a certain high school. The high school offers two Advanced Placement math / stat classes to seniors only: AP Calculus and AP Statistics. The roster of the Calculus class shows 95 people; the roster of the Statistics class shows 86 people. There are 43 overachieving seniors on both rosters.
12. What is the probability that a randomly selected senior is in at least one of the two classes?
(10 pts)

____________

13. If the student is in the Calculus class, what is the probability the student is also in the Statistics
class? (10 pts)

Answer by stanbon(75887)

There are 1000 students in the senior class at a certain high school. The high school offers two Advanced Placement math / stat classes to seniors only: AP Calculus and AP Statistics. The roster of the Calculus class shows 95 people; the roster of the Statistics class shows 86 people. There are 43 overachieving seniors on both rosters.
12. What is the probability that a randomly selected senior is in at least one of the two classes?
====
P(calc or stat) = [95+86-43]/1000 = 0.138
____________
13. If the student is in the Calculus class, what is the probability the student is also in the Statistics
P(stat | calc) = P(stat AND calc)/P(calc) = (43/1000)/(95/100) = 43/95
=====================================
Cheers,
Stan H.
=======================